2x+4y=8 and x+2y=4 the solution set of a pair of linear equation is
 

Grade:10

1 Answers

Shraddha Singh
2 years ago
Dear Student,
 
2x+4y = 8    --(i)
x + 2y = 4    --(ii)
 
Using substitution method, 
From eq(ii), 
X = 4-2y
 
Substituting value of x in eq(i)
2(4-2y) + 4y = 8
8-4y+4y = 8
8=8 
 
It means that it is an equation of coincident lines .
The solution set of this pair of linear equation is infinitely many solutions.
 
Hope it helps!!
Thank you
 
 
Using